# Batched Nonparametric Contextual Bandits

### v3

@@ 3 Fundamental limits of batched nonparametric bandits
In this section, we establish minimax lower bounds for the regret achievable by any $M$-batch policy $(\Gamma,\pi)$; see Theorem 2. To begin with, we state a minimax lower bound, together with its proof, when the grid $\Gamma$ is prespecified, that is, the statistician divides the horizon $[1:T]$ into $M$ disjoint batches $[1:t_{1}]$, $[t_{1}+1:t_{2}]$, $\ldots,[t_{M-1}+1,T]$ before the game begins; see Theorem 1. As we will soon see, the proof of the lower bound with fixed grid is not only useful for establishing the lower bound for any general $M$-batch policy $(\Gamma,\pi)$, but also instrumental in our development of an optimal policy to be detailed in Section 4.
@@ 3.1 Proof of Theorem 1
Our construction of the family of hard instances is adapted from [50]. Define $z_{1}=1$, and $z_{i}=\lceil t_{i-1}{}^{1/(2\beta+d)}\rceil$ for $i=2,3,\ldots,M$. Henceforth, we will fix some $i$ and write $z_{i}$ as $z$. We partition $[0,1]^{d}$ into $z^{d}$ bins with equal width. Denote the bins by $C_{j}$ for $j=1,...,z^{d}$, and let $q_{j}$ be the center of $C_{j}$.
@@ 3.2 Lower bound for general $M$-batch policy
$R_{T}(\pi)\geq\tilde{D}_{1}(\frac{1}{M})^{\tilde{D}_{2}}T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@
Unlike the fixed grid case where we choose a specific family of hard instances to target the regret in a certain batch, we cannot directly do so when the grid is adaptively selected because the adversary does not know $\{t_{i}\}_{i=1}^{M}$ in advance. Inspired by [21], we overcome this difficulty by using an appropriately defined bad event that happens with sufficient probability to reduce the adaptive case to the fixed grid case. However, the proof under the nonparametric setting is much more challenging because the presence of contexts makes the instances for different batches less indistinguishable with each other. A key ingredient of our proof is to establish tight control of the total variation distance between two mixture distributions. The full proof can be found in Appendix A.
@@ 4 Batched successive elimination with dynamic binning
        • Update $\mathcal{L}$ and $\{\mathcal{I}_{C}\}_{C\in\mathcal{L}}$ by Algorithm 2 (,$(\mathcal{L},\{\mathcal{I}_{C}\}_{C\in\mathcal{L}},i,g_{i})$.
@@
Figure 1: An example of the tree growing process for $d=1,M=3,G=\{4,3,1\}$. The root node is at depth 0. For the first batch, the 4 nodes located at depth 1 of the tree were used. Both $[\frac{1}{4},\frac{1}{2})$ and $[\frac{1}{2},\frac{3}{4})$ only had one active arm remaining so they were not further split and remained in the set of active nodes (green). Meanwhile, $|\mathcal{I}_{[0,\frac{1}{4})}|=|\mathcal{I}_{[\frac{3}{4},1]}|=2$ so each of them was split into 3 smaller nodes, and both nodes were marked as inactive (red). For the second batch, all the green nodes were actively used but arm elimination was performed at the end of batch 2 only for nodes located at depth 2 (the green nodes at depth 1 already have 1 active arm remaining so there is no need to eliminate again).
@@ 4.3 Failure of static binning
Consider $M=3$, and $\alpha=\beta=d=1$. For any choice of $11\leq t_{1}<t_{2}\leq T-$ 1, and any choice of $g$, there exists a nonparametric bandit instance in $\mathcal{F}(1,1)$ such that the resulting successive elimination with static binning algorithm $\hat{\pi}_{\mathrm{static}}$ satisfies
@@ Appendix A Proof of Theorem 2
It is worth emphasizing that Theorem 2 aims to establish the hardness of batched nonparametric bandits even when the grid $\Gamma$ is allowed to be adaptively chosen. Nevertheless, the proof of Theorem 1 in the fixed-grid case is still useful.
We reuse the family of hard instances $\mathcal{C}_{z_{i}}$ as defined in (4), and define the mixture distribution
$Q_{i}(\cdot)=\frac{1}{s_{i}2^{s_{i}-1}}\sum_{j=1}^{s_{i}}\sum_{\omega_{[-j]}\in\Omega_{s_{i}-1}}\mathbb{P}_{\pi,i,f_{\omega_{[-j]}^{-1}}}(\cdot)=\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)\mathbb{P}_{\pi,i,f_{\omega}}(\cdot),$ (13)
where $q_{i}:\Omega_{s_{i}}\rightarrow[0,1]$ is selected so that the above equality holds, and $\mathbb{P}_{\pi,i,f_{\omega}}$ is the distribution of the observations when $f_{\omega}\in\mathcal{C}_{z_{i}}$. It is easy to see $\sum_{\omega\in\Omega_{s_{i}}}q_{i}(\omega)=1$. Let $\{T_{i}\}_{i=1}^{M}$ and $\{z_{i}\}_{i=1}^{M}$ be strictly increasing sequences of positive integers to be specified later. We pause here to state a useful claim regarding the family $\{Q_{i}\}_{i=1}^{M}$ of mixture distributions, namely, they are all close to each other under the total variation distance.
Lemma 2.
For any $1\leq i\leq M$, one has $\mathrm{TV}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})\leq\frac{1}{2}\sqrt{T_{i-1}z_{i}^{-(2\beta+d)}}.$
The following lemma ensures that when a sequence of events $A_{i}$’s are properly designed, at least one of them happens with sufficiently large probability under the mixture distribution.
Lemma 3.
Assume $z_{i}=\lceil(36T_{i-1}(M-k+1)^{2})^{1/(2\beta+d)}\rceil$. Suppose $\{A_{i}\}_{i=k}^{M}$ is a sequence of events that partition the whole probability space where $1\leq k\leq M$. Assume that each $A_{i}$ can be determined by observations up to $T_{i-1}$. Then, there exists $i^{\star}\in\{k,k+1,...,M\}$ such that $Q_{i^{\star}}(A_{i^{\star}})\geq 1/(2(M-k+1))$.
The next lemma indeed shows that when $A_{i}$ happens, the regret must be large.
Lemma 4.
Let $A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}$. If $Q_{i}(A_{i})\geq 1/(2(M-k+1))$, then
$\sup_{f\in\mathcal{C}_{z_{i}}}R_{T_{i}}(\pi;f)\apprge T_{i}z_{i}^{-\beta(1+\alpha)}(M-k+1)^{-\frac{1+\alpha}{\alpha}}.$
@@
The proofs of the above lemmas are postponed to the end of this section. Now we are ready to establish the desired claim in the theorem. Define $b\asymp T^{(1-\gamma)/(1-\gamma^{M})}$. For each $1\leq i\leq M$, we set $T_{i}=\lfloor b^{(1-\gamma^{i})/(1-\gamma)}\rfloor$, $z_{i}=\lceil(36T_{i-1}M^{2})^{1/(2\beta+d)}\rceil$, and $s_{i}\coloneqq\lceil z_{i}^{d-\alpha\beta}\rceil$. Define the event
$A_{i}=\{t_{i-1}<T_{i-1}<T_{i}\leq t_{i}\}.$
@@
Intuitively, $A_{i}$ models the event when the algorithm’s selected grid points $t_{i-1}$ and $t_{i}$ are suboptimal. When $A_{i}$ happens, the goal is to design a problem instance such that using observations up $t_{i-1}$ cannot distinguish the optimal arm and therefore the policy must incur a large regret between $t_{i-1}$ and $t_{i}$. Such $A_{i}$’s satisfy the assumption in Lemma 3 with $k=1$ and we can combine Lemma 3 and Lemma 4 to obtain
$\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha,\beta)}R_{T}(\pi;f)\geq\sup_{f\in\mathcal{C}_{z_{i^{\star}}}}R_{T_{i^{\star}}}(\pi;f)\apprge T_{i^{\star}}z_{i^{\star}}^{-\beta(1+\alpha)}M^{-\frac{1+\alpha}{\alpha}}=\tilde{D}_{1}(\frac{1}{M})^{\tilde{D}_{2}}T^{\frac{1-\gamma}{1-\gamma^{M}}},$
where the second inequality uses Lemma 4, and the last one arises from the definitions of $T_{i}$ and $z_{i}$.
A.1 Proof of Lemma 2
It suffices to bound their KL divergence. By the standard decomposition of KL divergence and Bernoulli reward structure,
@@ A.1 Proof of Lemma 2
$\mathrm{KL}(Q_{M}^{T_{i-1}},Q_{i}^{T_{i-1}})$ $\leq 8\sum_{t=1}^{T_{i-1}}\mathbb{E}_{Q_{M}}\left[\left(\underbrace{\sum_{\omega\in\Omega_{M}}q_{M}(\omega)f_{M,\omega}(X_{t})-\sum_{\omega\in\Omega_{i}}q_{i}(\omega)f_{i,\omega}(X_{t})}_{\Delta_{t}}\right)^{2}\mathbf{1}\{\pi_{t}(X_{t})=1\}\right],$ (14)
where $f_{i,\omega}$ denotes an instance from $\mathcal{C}_{z_{i}}$. To control $\Delta_{t}$, we further decompose it as
@@
$\Delta_{t}$ $=\sum_{j=1}^{s_{i}}\left(\sum_{\omega\in\Omega_{M}}q_{M}(\omega)f_{M,\omega}(X_{t})-\sum_{\omega\in\Omega_{i}}q_{i}(\omega)f_{i,\omega}(X_{t})\right)\mathbf{1}\{X_{t}\in C_{i,j}\},$
where $C_{i,j}$ is the $j^{th}$ bin corresponding to the instance family $\mathcal{C}_{z_{i}}$. Here, the difference between the two sums can be restricted to $\cup_{j=1}^{s_{i}}C_{i,j}$ because $\cup_{j=1}^{s_{M}}C_{M,j}\subseteq\cup_{j=1}^{s_{i}}C_{i,j}$. Indeed, the area of effective bins for an instance in $\mathcal{C}_{z_{i}}$ is $s_{i}z_{i}^{-d}=z_{i}^{-\alpha\beta}$, which decreases as $i$ increases. Notice for $X_{t}\in C_{i,j}$,
$|\sum_{\omega\in\Omega_{i}}q_{i}(\omega)f_{i,\omega}(X_{t})-\frac{1}{2}|\leq\frac{2^{s_{i}-1}}{s_{i}\cdot 2^{s_{i}-1}}\cdot\frac{z_{i}^{-\beta}}{4}=\frac{z_{i}^{-\beta}}{4s_{i}},$
because all the $\omega_{[-j]}^{-1}$’s have a negative sign in the $j^{th}$ bin and there are $2^{s_{i}-1}$ of them, while for each $k\neq j$, the positive and negative spikes within the $j^{th}$ bin cancel out each other when summing over $\omega_{[-k]}^{-1}$’s due to symmetry. Therefore,
$|\Delta_{t}|\leq\frac{z_{i}^{-\beta}}{4s_{i}}\sum_{j=1}^{s_{i}}\mathbf{1}\{X_{t}\in C_{i,j}\}=\frac{1}{4}z_{i}^{-\beta-d+\alpha\beta}\sum_{j=1}^{s_{i}}\mathbf{1}\{X_{t}\in C_{i,j}\}.$
@@
Plugging the above back to (14) we obtain

### v4

@@ 3 Fundamental limits of batched nonparametric bandits
[…]
@@ 3.1 Proof of Theorem 1
Our construction of the family of hard instances is adapted from [50]. Define $z_{1}=1$, and $z_{i}=\lceil t_{i-1}{}^{1/(2\beta+d)}\rceil$ for $i=2,3,\ldots,M$. Henceforth, we will fix some $i$ and write $z_{i}$ as $z$. We partition $0,1]^{d}$ into $z^{d}$ bins with equal width. Denote the bins by $C_{j}$ for $j=1,...,z^{d}$, and let $q_{j}$ be the center of $C_{j}$.
@@ 3.2 Lower bound for general $M$-batch policy
$R_{T}(\pi)\geq\tilde{D}_{1}(\frac{1}{M})^{\tilde{D}_{2}}\cdot T^{\frac{1-\gamma}{1-\gamma^{M}}},$
@@
Unlike the fixed grid case where we choose a specific family of hard instances to target the regret in a certain batch, we cannot directly do so when the grid is adaptively selected because the adversary does not know $\{t_{i}\}_{i=1}^{M}$ in advance. Inspired by [21], we overcome this difficulty by using an appropriately defined bad event that happens with sufficient probability to reduce the adaptive case to the fixed grid case. A significant technical difference to MAB lies in the construction of a family of nonparametric bandit instances that are simultaneously hard for all batches. The full proof can be found in Appendix A.
@@ 4 Batched successive elimination with dynamic binning
        • Update $\mathcal{L}$ and $\{\mathcal{I}_{C}\}_{C\in\mathcal{L}}$ by Algorithm 2 $(\mathcal{L},\{\mathcal{I}_{C}\}_{C\in\mathcal{L}},i,g_{i})$.
@@
Figure 1: An example of the tree growing process for $d=1,M=3,G=\{4,3,1\}$. The root node is at depth 0. For the first batch, the 4 nodes located at depth 1 of the tree were used. Both $\frac{1}{4},\frac{1}{2})$ and $\frac{1}{2},\frac{3}{4})$ only had one active arm remaining so they were not further split and remained in the set of active nodes (green). Meanwhile, $|\mathcal{I}_{[0,\frac{1}{4})}|=|\mathcal{I}_{[\frac{3}{4},1]}|=2$ so each of them was split into 3 smaller nodes, and both nodes were marked as inactive (red). For the second batch, all the green nodes were actively used but arm elimination was performed at the end of batch 2 only for nodes located at depth 2 (the green nodes at depth 1 already have 1 active arm remaining so there is no need to eliminate again).
@@ 4.3 Failure of static binning
Consider $M=3$, and $\alpha=\beta=d=1$. For any choice of $1\leq t_{1}<t_{2}\leq T-1$, and any choice of $g$, there exists a nonparametric bandit instance in $\mathcal{F}(1,1)$ such that the resulting successive elimination with static binning algorithm $\hat{\pi}_{\mathrm{static}}$ satisfies
@@ Appendix A Proof of Theorem 2
Define $b\asymp T^{(1-\gamma)/(1-\gamma^{M})}$. For each $1\leq m\leq M$, we set $T_{m}=\lfloor b^{(1-\gamma^{m})/(1-\gamma)}\rfloor$, $z_{m}=\lceil(36T_{m-1}M^{2})^{1/(2\beta+d)}\rceil$. Define the event
$A_{m}=\{t_{m-1}<T_{m-1}<T_{m}\leq t_{m}\}.$
@@
Intuitively, $A_{m}$ models the event when the algorithm’s selected grid points $t_{m-1}$ and $t_{m}$ are suboptimal. When $A_{m}$ happens, the goal is to design a problem instance such that using observations up $t_{m-1}$ cannot distinguish the optimal arm and therefore the policy must incur a large regret between $t_{m-1}$ and $t_{m}$.
A.1 Construction of the hard instances
Consider a regular partition of $0,1]^{d}$ into $M^{d}$ bins of equal width. Denote the bins by $C_{i}$ for $i=1,\dots,M^{d}$. For any $m\in[M],$ define $I_{m}=\{i\in[M^{d}]:(m-1)M^{d-1}+1\leq i\leq mM^{d-1}\}$. Given $i\in I_{m}$, we further consider a regular partition of $C_{i}$ into $z_{m}^{d}$ bins, which are denoted by $C_{i,j}^{m}$ for $j=1,\dots,z_{m}^{d}$. Denote by $q_{i,j}^{m}$ the center of $C_{i,j}^{m}$.
Define $s_{m}=\lceil(M^{-\alpha\beta})z_{m}^{d-\alpha\beta}\rceil$ and $s=M^{d-1}\sum_{m=1}^{M}s_{m}$. Denote by $\Omega_{m}=\{\pm 1\}^{s_{m}}$. Define a set of binary sequences
$\Omega=\{\omega\in\{\pm 1\}^{s}:1\leq m\leq M,i\in I_{m},\omega_{i}\in\Omega_{m}\},$
where for $i\in I_{m}$, $\omega_{i}$ stands for the $i$-th block in $\omega\in\{\pm 1\}^{s}$ of size $s_{m}$. For any $\omega\in\Omega$, define a function $f_{\omega}:[0,1]^{d}\mapsto\mathbb{R}$:
$f_{\omega}(x)=\frac{1}{2}+\sum_{m=1}^{M}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}\omega_{i,j}^{m}\xi_{i,j}^{m}(x),$
where $\xi_{i,j}^{m}(x)=D_{\phi}(Mz_{m})^{-\beta}\phi(2Mz_{m}(x-q_{i,j}^{m}))\mathbf{1}\{x\in C_{i,j}^{m}\}$. It is straightforward to check $f_{\omega}$ satisfies the smoothness condition.
Define the family of reward instances
$\mathcal{C}\coloneqq\left\{f_{1}(x)=f_{\omega}(x),f_{-1}(x)=\frac{1}{2}\mid\omega\in\Omega\right\}.$
We now verify the margin condition. One has
$P_{X}\left(0<\left|f_{\omega}(X)-\frac{1}{2}\right|\leq D_{\phi}\delta\right)$ $=\sum_{m=1}^{M}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}P_{X}(0<\left|f_{\omega}(X)-\frac{1}{2}\right|\leq D_{\phi}\delta,X\in C_{i,j}^{m})$
$=\sum_{m=1}^{M}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}P_{X}(0<\phi(2Mz_{m}(x-q_{i,j}^{m}))\leq\delta(Mz_{m})^{\beta},X\in C_{i,j}^{m})$
$\leq\sum_{m=1}^{M}\sum_{i\in I_{m}}M^{-d}(1+d)\delta^{\alpha}=(1+d)\delta^{\alpha},$
where the penultimate step applies Lemma 3. Hence, the margin assumption is satisfied, and $\mathcal{C}\subseteq\mathcal{F}(\alpha,\beta)$.
A.2 Lower bounding the regret during the $m$-th batch
Since the worst-case inferior sampling rate is lower bounded by the average over the family $\Omega$,
$\sup_{(f,\frac{1}{2})\in\mathcal{F}(\alpha,\beta)}$ $S_{T_{m}}(\pi,f)$
@@ A.2 Lower bounding the regret during the $m$-th batch
$\geq\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{E}_{\pi,\omega}\left[\sum_{t=1}^{T_{m}}\mathbf{1}\{\pi_{t}(X_{t})\neq\pi^{\star}(X_{t}),f_{\omega}(X_{t})\neq\frac{1}{2}\}\right]$
@@
$\overset{\mathrm{(i)}}{\geq}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{E}_{\pi,\omega}^{t}\left[\mathbf{1}\{X_{t}\in C_{i,j}^{m},\pi_{t}(X_{t})\neq\omega_{i,j}^{m}\}\right]$
$=(Mz_{m})^{-d}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}\frac{1}{2^{s}}\sum_{\omega_{-(i,j)}\in\Omega_{-(i,j)}}\underbrace{\sum_{l\in\{\pm 1\}}\mathbb{E}_{\pi,\omega_{i,j}^{m}=l}^{t}P_{X}(\pi_{t}(X_{t})\neq l\mid X_{t}\in C_{i,j}^{m})}_{U_{m,i,j}^{t}}.$ (13)
Here, step (i) uses the fact that regret is only incurred on $C_{i,j}^{m}$’s and the optimal action is specified by $\omega_{i,j}^{m}$. Besides, we use $\omega_{-(i,j)}$ to represent the vector after leaving out the $j$-th entry in the $i$-th block of $\omega$. By Le Cam’s method, one has
$U_{m,i,j}^{t}$ $\geq 1-\|\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{t}-\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{t}\|_{\mathrm{TV}}$
$\geq 1-\|\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m}}-\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m}}\|_{\mathrm{TV}}$
$=\int\min\left\{\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m}},\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m}}\right\}$
$\geq\int_{A_{m}}\min\left\{\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m}},\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m}}\right\},$
where the second inequality is due to $t\leq T_{m}$. Since the available observations for $\pi$ at $T_{m}$ are the same as those at $T_{m-1}$ under $A_{i}$, we continue to lower bound
$U_{m,i,j}^{t}$ $\geq\int_{A_{m}}\min\left\{\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m-1}},\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m-1}}\right\}$
$=\frac{1}{2}\int_{A_{m}}\left(\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m-1}}+\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m-1}}-|\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m-1}}-\mathrm{d}\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m-1}}|\right)$
$\geq\frac{1}{2}\left(\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m-1}}(A_{m})+\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m-1}}(A_{m})\right)-\|\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}^{T_{m-1}}-\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}^{T_{m-1}}\|_{\mathrm{TV}}$
$\geq\frac{1}{2}\left(\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}(A_{m})+\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}(A_{m})\right)-\frac{1}{2M},$
@@
where the last step applies Lemma 4. Plugging the above back to (13), we obtain
$\sup_{f\in\mathcal{F}(\alpha,\beta)}$ $S_{T_{m}}(\pi,f)$
$\geq(Mz_{m})^{-d}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}\frac{1}{2^{s+1}}\sum_{\omega_{-(i,j)}\in\Omega_{-(i,j)}}\left(\mathbb{P}_{\pi,\omega_{i,j}^{m}=-1}(A_{m})+\mathbb{P}_{\pi,\omega_{i,j}^{m}=1}(A_{m})-\frac{1}{M}\right)$
$=(Mz_{m})^{-d}\sum_{t=T_{m-1}+1}^{T_{m}}\sum_{i\in I_{m}}\sum_{j=1}^{s_{m}}\frac{1}{2}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2M}\right)$
$=(Mz_{m})^{-d}(T_{m}-T_{m-1})M^{d-1}s_{m}\left(\mathbb{E}_{\omega\sim\mathrm{Unif}(\Omega)}\mathbb{P}_{\pi,\omega}(A_{m})-\frac{1}{2M}\right)$